# Hodge diamonds of the Landau–Ginzburg orbifolds

### 26 Feb 2024 v2

@@ 4.5. Bigrading
$(q_{l},q_{r})=\left(\deg p-\sum_{k\in I_{g}^{c}}q_{k}+\mathrm{age}{(g)},\deg p-\sum_{k\in I_{g}^{c}}q_{k}+\mathrm{age}{(g^{-1})}\right).$ (39)
This definition endows ${\mathcal{B}}_{tot}$ with the structure of a ${\mathbb{Q}}$-bigraded vector space. For $u,v\in G_{f}^{d}$ it follows immediately that
$q_{\bullet}(\xi_{u})+q_{\bullet}(\xi_{v})=q_{\bullet}(\xi_{uv})$
for $u,v\in G$, such that $I_{u}^{c}\cap I_{v}^{c}=\emptyset$.
This bigrading restricts to ${\mathcal{B}}(f,G)$ because $q_{l},q_{r}$ commute with the action of $h^{\ast}$ for any $h\in{\mathrm{GL}}_{f}$, $h$ preserves the weights and $\mathrm{age}(g)=\mathrm{age}(hgh^{-1})$.
5. Hodge diamond of LG orbifolds
@@
$(j_{f})^{*}([p]\xi_{g})={\bf e}[-\deg(p)+\sum_{k\in I_{g}^{c}}q_{k}]\cdot[p]\xi_{g}$
by Example 31. Hence for a class in ${\mathcal{B}}(f,G)$ we have ${\bf e}[-q_{l}+\mathrm{age}(g)]=1$ and so $q_{l}$ is integer. ∎
@@
Proposition 35.
Let $f$ be a quasihomogeneous polynomial satsifying the CY condition, let $G\subseteq{\mathrm{SL}}_{f}$ be a finite subgroup, and let $V^{a,b}$ stand for the bidegree $(a,b)$–subspace of ${\mathcal{B}}(f,G)$. We have
@@
$q_{l}(X)=\deg p-\sum_{k\in I_{g}^{c}}q_{k}+\mathrm{age}(g)$
we see that $q_{l}(X)\geq 0$ because $\sum_{k\in I_{g}^{c}}q_{k}\leq\sum_{k=1}^{N}q_{k}=1$. Similarly for $q_{r}(X)$ by the same argument applied to $\mathrm{age}(g^{-1})$.
(ii) If $g=\mathrm{id}$ we have that $q_{l}(X)=q_{r}(X)=0$ if and only if $\deg p=0$. By Propositions 28 and 29 we have $[p]=\alpha[1]$ in $\mathrm{Jac}(f)$ for some constant $\alpha\in{\mathbb{C}}$.
For $g\neq\mathrm{id}$ we just saw that $\mathrm{age}(g)\geq 1$ and $\sum_{k\in I_{g}^{c}}q_{k}\leq 1$, so $q_{l}(X)=q_{r}(X)=0$ is achieved only if $\deg p=0$, $N_{g}=0$ and $\mathrm{age}(g)=\mathrm{age}(g^{-1})=1$, which implies $N=\mathrm{age}(g)+\mathrm{age}(g^{-1})+N_{g}=2$.
(iii) If $g=\mathrm{id}$, the statement follows from Proposition 28 as given the CY condition we have $\hat{c}=N-2\sum_{k}q_{k}=N-2$.
For $g\neq\mathrm{id}$ apply the same proposition again to estimate $\deg p$ in $\mathrm{Jac}(f^{g})$. Namely, it gives
@@
At the same time we have $N_{g}+\mathrm{age}(g)=N-\mathrm{age}(g^{-1})\leq N-1$ because $\mathrm{age}(g^{-1})\in{\mathbb{N}}_{\geq 1}$. Combining this with the inequality above we get
$q_{l}(X)\leq N-2-\sum_{k\in I_{g}}q_{k}\leq N-2.$ (40)
@@
(iv) If $g=\mathrm{id}$, by Proposition 28 we see that $q_{l}([\mathrm{hess}(f)]\xi_{\mathrm{id}})=q_{r}([\mathrm{hess}(f)]\xi_{\mathrm{id}})=N-2$.
If $g\neq\mathrm{id}$, $q_{l}(X)=q_{r}(X)=N-2$, then Eq. (40) implies that $\sum_{k\in I_{g}}q_{k}=0$ and, hence, $N_{g}=0$ and $\mathrm{age}(g)=\mathrm{age}(g^{-1})=1$, which altogether implies $N=\mathrm{age}(g)+\mathrm{age}(g^{-1})+N_{g}=2$.
We now construct two symmetries of ${\mathcal{B}}_{tot}$. The horizontal morphism $\Psi$ and the vertical morphism $\Phi$.
Consider the direct sum decomposition of ${\mathcal{B}}_{tot}$ as in Eq. (30). We first extend the $\phi_{f,\beta}$ isomorphism of Proposition 29 to ${\mathcal{B}}_{tot}$ in the following way. The hessian matrix of $f^{g}$ viewed coordinate free is a bilinear form on the tangent bundle of $\mathrm{Fix}(g)$. Therefore, its determinant $\mathrm{hess}(f^{g})$ is canonically an element of $(\Lambda^{N_{g}}\mathrm{Fix}(g)^{\vee})^{\otimes 2}\otimes\mathbb{C}[\mathrm{Fix}(g)]$. Fix a generator $\xi^{\vee}_{g}:=\iota_{\xi_{g}}dx_{1}\wedge\ldots\wedge dx_{N}\in\Lambda^{N_{g}}\mathrm{Fix}(g)^{\vee}$, where $\iota$ is the interior product operator and let the generator of $(\Lambda^{N_{g}}\mathrm{Fix}(g)^{\vee})^{\otimes 2}$ to be $(\xi^{\vee}_{g})^{\otimes 2}$. This choice allows us to fix $\mathrm{hess}(f^{g})$ as a function on $\mathrm{Fix}(g)$ and, hence fix a pairing $\eta_{f^{g}}$ for the $g$-sector. As in Proposition 29 this in turn defines an isomorphism $\phi_{f^{g},\beta}$ on each sector. Now the vertical morphism $\Phi$ is the direct sum of these isomorphisms acting on each sector of ${\mathcal{B}}_{tot}$
$\Phi:=\bigoplus_{g\in{\mathrm{GL}}_{f},\beta\in{\mathbb{Q}}}\phi_{f^{g},\beta}\ \colon\ {\mathcal{B}}_{tot}\to{\mathcal{B}}_{tot}^{\vee}.$
It is an isomorphism restricted to ${\mathcal{B}}_{tot,G}$ for any finite $G$ because each of $\phi_{f^{g},\beta}$ is an isomorphism.
Define the horizontal morphism $\Psi$ to act on the $g$–th sector by
$\Psi([p]\xi_{g}):=[p]\xi_{g^{-1}}.$
Extend it by linearity to all ${\mathcal{B}}_{tot}$. This is an isomorphism because $f^{g}=f^{g^{-1}}$ and $\mathrm{Jac}(f^{g})=\mathrm{Jac}(f^{g^{-1}})$.
Proposition 36.
1) The maps $\Phi$ and $\Psi$ are well defined on ${\mathcal{B}}(f,G)$ for any finite $G\subseteq{\mathrm{SL}}_{f}$.
2) For $f$ satisfying CY condition and a finite $G\subseteq{\mathrm{SL}}_{f}$, let $V^{a,b}$ stand for the bidegree $(a,b)$–subspace of ${\mathcal{B}}(f,G)$. Then the maps $\Psi$ and $\Phi$ induce the ${\mathbb{C}}$–vector spaces isomorphisms
$V^{a,b}\cong V^{b,a}\quad\text{ and }\quad V^{a,b}\cong(V^{N-2-b,N-2-a})^{\vee}.$
@@
1) The map $\Psi$ commutes with the $G$-action since $\mathrm{Fix}(g)=\mathrm{Fix}(g^{-1})$. Hence $\Psi$ preserves the invariants.
To see that $\Phi$ commute with the $G$-action, recall first that, $f(h^{-1}\cdot\mathbf{x})=f(\mathbf{x})$ and, hence, $f^{g}(h^{-1}\cdot\mathbf{x})=f^{hgh^{-1}}(\mathbf{x})$. Furthermore, since $\det(h)=1$ we have $\rho_{h,g}h^{*}(\xi_{hgh^{-1}}^{\vee})=\xi^{\vee}_{g}$ and we can conclude that
$\rho_{h,g}^{2}\eta_{f^{hgh^{-1}}}([p_{1}(h^{-1}\cdot\mathbf{x})]\xi_{hgh^{-1}},[p_{2}(h^{-1}\cdot\mathbf{x})]\xi_{hgh^{-1}})=\eta_{f^{g}}([p_{1}(\mathbf{x})]\xi_{g},[p_{2}(\mathbf{x})]\xi_{g}).$
@@
To verify compatibility of $\Phi$ with the grading, note first, that $\widehat{c}(f^{g})=\sum_{k\in I_{g}}(1-2q_{k})$. Thus, by Proposition 29 the left charge of $[\phi_{f^{g},\deg p}(p)]\xi_{g}$ is given by
$q_{l}([\phi_{f^{g},\deg p}(p)]\xi_{g})$ $=\sum_{k\in I_{g}}(1-2q_{k})-\deg p-\sum_{k\in I_{g}^{c}}q_{k}+\mathrm{age}{(g)}=$
$=N_{g}-2\sum_{k\in I_{g}}q_{k}-\deg p-\sum_{k\in I_{g}^{c}}q_{k}+(N-N_{g}-\mathrm{age}{(g^{-1})})$
$=-2+\sum_{k\in I^{c}_{g}}q_{k}-\deg p+N-\mathrm{age}{(g^{-1})}=N-2-q_{r}([p]\xi_{g}).$
The computation for the right charge is identical.
@@
Proposition 37.
For $f$ satisfying CY condition and a finite $G$, such that $J\subseteq G\subseteq{\mathrm{SL}}_{f}$, let $V^{a,b}$ stand for the bidegree $(a,b)$–subspace of ${\mathcal{B}}(f,G)$. Then
@@
One notes immediately that $[1]\xi_{j_{f}}$ and $[1]\xi_{j_{f}^{-1}}$ are non-zero in ${\mathcal{B}}(f,G)$ and belong to $V^{0,N-2}$ and $V^{N-2,0}$ respectively. By Proposition 36 it’s enough to show one of the statements.
Lemma 38.
$\dim V^{0,N-2}=\{g\in G\ |\ \mathrm{age}(g)=1,\ N_{g}=0\}$.
@@
Let $[p]\xi_{g}\in V^{0,N-2}$. It follows from Eq. (14) and (39) that $\mathrm{age}(g)=1-N_{g}/2$. The statement follows by Proposition 19. ∎
Under the CY condition for $g\in{\mathrm{GL}}_{f}\backslash\{\mathrm{id}\}$ of finite order and with integral $\mathrm{age}(g)$ we have by Proposition 26 that $\mathrm{age}(g)\geq 1$ with the equality being reached only for $g=j_{f}$. This completes the proof. ∎
This completes the proof of Theorem 1.
@@
$h^{a,b}:=\dim_{\mathbb{C}}\left\{X\in{\mathcal{B}}(f,G)\ |\ (q_{l}(X),q_{r}(X))=(a,b)\right\}$ (41)
@@
It follows from the propositions above that for $f$ satisfying CY condition and $G$ such that $J\subseteq G\subseteq{\mathrm{SL}}_{f}$, the numbers $h^{a,b}$ form a diamond.
{tikzpicture}
Let’s call the line $\{h^{a,b}\ |\ a+b=D\}$ the horizontal line and the line $\{h^{a,b}\ |\ a-b=0\}$ the vertical line. The Hodge diamond $\{h^{a,b}\}_{a,b=0}^{D}$ has the following special properties
@@
    The $g$–th sector of ${\mathcal{B}}(f,G)$ contributes as a line symmetric with respect to the horizontal line.
@@
    Every $g$–th sector of ${\mathcal{B}}(f,G)$ contributes together with a $g^{-1}$–th sector of ${\mathcal{B}}(f,G)$, located symmetrically with respect to the vertical line.
@@
$\xi_{j_{f}^{3}},\ \xi_{j_{f}^{5}},\ \xi_{j_{f}^{7}},\ \xi_{j_{f}^{9}},\ [x_{1}]\xi_{j_{f}^{4}},\ [x_{1}]\xi_{j_{f}^{8}},\ [x_{4}^{2}]\xi_{j_{f}^{6}},$ (42)
$[x_{3}^{4}x_{4}^{4}]\xi_{\mathrm{id}},\ [x_{1}x_{3}x_{4}^{4}]\xi_{\mathrm{id}},\ [x_{1}x_{3}^{3}x_{4}^{3}]\xi_{\mathrm{id}},\ [x_{2}x_{4}^{3}]\xi_{\mathrm{id}},\ [x_{1}^{2}x_{4}^{3}]\xi_{\mathrm{id}},\ [x_{1}x_{3}^{5}x_{4}^{2}]\xi_{\mathrm{id}},\ [x_{2}x_{3}^{2}x_{4}^{2}]\xi_{\mathrm{id}},$ (43)
$[x_{1}^{2}x_{3}^{2}x_{4}^{2}]\xi_{\mathrm{id}},\ [x_{2}x_{3}^{4}x_{4}]\xi_{\mathrm{id}},\ [x_{1}x_{2}x_{3}x_{4}]\xi_{\mathrm{id}},\ [x_{1}^{3}x_{3}x_{4}]\xi_{\mathrm{id}},\ [x_{1}^{2}]\xi_{\mathrm{id}},\ [x_{2}^{2}]\xi_{\mathrm{id}}.$ (44)
@@
$\xi_{j_{f}},\quad\xi_{j_{f}^{11}},\quad[1]\xi_{\mathrm{id}},\quad[x_{1}x_{2}^{2}x_{3}x_{4}^{4}]\xi_{\mathrm{id}},$
having the bigrading $(0,2)$, $(2,0)$, $(0,0)$ and $(2,2)$ respectively.
One gets the following diamond
{tikzpicture}
Example 40.
@@
$\xi_{\mathrm{id}}\enskip\text{and}\enskip[x_{1}^{3}x_{2}^{3}x_{3}^{3}x_{4}^{3}x_{5}^{3}]\xi_{\mathrm{id}}$
in bidegrees $(0,0)$ and $(3,3)$ respectively;
$[x_{1}x_{2}x_{3}x_{4}x_{5}]\xi_{\mathrm{id}},[x_{4}^{3}x_{5}^{2}]\xi_{\mathrm{id}},[x_{4}^{2}x_{5}^{3}]\xi_{\mathrm{id}},[x_{1}x_{2}x_{3}x_{4}^{2}]\xi_{\mathrm{id}},[x_{1}x_{2}x_{3}x_{5}^{2}]\xi_{\mathrm{id}},$
$[(x_{1}+x_{2}+x_{3})^{2}](\xi_{(1,2,3)}+\xi_{(1,3,2)}),[x_{4}^{2}](\xi_{(1,2,3)}+\xi_{(1,3,2)}),[x_{5}^{2}](\xi_{(1,2,3)}+\xi_{(1,3,2)}),$
$[(x_{1}+x_{2}+x_{3})x_{4}](\xi_{(1,2,3)}+\xi_{(1,3,2)}),[(x_{1}+x_{2}+x_{3})x_{5}](\xi_{(1,2,3)}+\xi_{(1,3,2)}),[x_{4}x_{5}](\xi_{(1,2,3)}+\xi_{(1,3,2)})$
@@
$\xi_{j_{f}}\enskip\text{and}\enskip\xi_{j_{f}^{4}}$
in bidegrees $(3,0)$ and $(0,3)$ respectively;
$\xi_{j_{f}^{2}},\xi_{(1,2,3)j_{f}}+\xi_{(1,3,2)j_{f}},\xi_{(1,2,3)j_{f}^{2}}+\xi_{(1,3,2)j_{f}^{2}}$

### 25 Mar 2024 v3

@@ 4.5 Bigrading
$(q_{l},q_{r})=\biggl{(}\deg p-\sum_{k\in I_{g}^{c}}q_{k}+\mathrm{age}{(g)},\deg p-\sum_{k\in I_{g}^{c}}q_{k}+\mathrm{age}{\bigl{(}g^{-1}\bigr{)}}\biggr{)}.$ (4.3)
This definition endows ${\mathcal{B}}_{\rm tot}$ with the structure of a ${\mathbb{Q}}$-bigraded vector space. For $u,v\in G_{f}^{d}$ it follows immediately that $q_{\bullet}(\xi_{u})+q_{\bullet}(\xi_{v})=q_{\bullet}(\xi_{uv})$ for $u,v\in G$, such that $I_{u}^{c}\cap I_{v}^{c}=\varnothing$.
This bigrading restricts to ${\mathcal{B}}(f,G)$ because $q_{l}$, $q_{r}$ commute with the action of $h^{\ast}$ for any $h\in{\mathrm{GL}}_{f}$, $h$ preserves the weights and $\mathrm{age}(g)=\mathrm{age}\bigl{(}hgh^{-1}\bigr{)}$.
5 Hodge diamond of LG orbifolds
@@
$(j_{f})^{*}([p]\xi_{g})={\bf e}\biggl{[}-\deg(p)+\sum_{k\in I_{g}^{c}}q_{k}\biggr{]}\cdot[p]\xi_{g}$
by Example 4.4. Hence for a class in ${\mathcal{B}}(f,G)$ we have ${\bf e}[-q_{l}+\mathrm{age}(g)]=1$ and so $q_{l}$ is integer. ∎
@@
Proposition 5.2.
Let $f$ be a quasihomogeneous polynomial satisfying the CY condition, let $G\subseteq{\mathrm{SL}}_{f}$ be a finite subgroup, and let $V^{a,b}$ stand for the bidegree $(a,b)$-subspace of ${\mathcal{B}}(f,G)$. We have
@@ 4. $(iv)$
$q_{l}(X)=\deg p-\sum_{k\in I_{g}^{c}}q_{k}+\mathrm{age}(g),$
we see that $q_{l}(X)\geq 0$ because $\sum_{k\in I_{g}^{c}}q_{k}\leq\sum_{k=1}^{N}q_{k}=1$. Similarly for $q_{r}(X)$ by the same argument applied to $\mathrm{age}\bigl{(}g^{-1}\bigr{)}$.
(ii) If $g=\mathrm{id}$ we have that $q_{l}(X)=q_{r}(X)=0$ if and only if $\deg p=0$. By Propositions 4.1 and 4.2, we have $[p]=\alpha[1]$ in $\mathrm{Jac}(f)$ for some constant $\alpha\in{\mathbb{C}}$.
For $g\neq\mathrm{id}$, we just saw that $\mathrm{age}(g)\geq 1$ and $\sum_{k\in I_{g}^{c}}q_{k}\leq 1$, so $q_{l}(X)=q_{r}(X)=0$ is achieved only if $\deg p=0$, $N_{g}=0$ and $\mathrm{age}(g)=\mathrm{age}(g^{-1})=1$, which implies
$N=\mathrm{age}(g)+\mathrm{age}\bigl{(}g^{-1}\bigr{)}+N_{g}=2.$
(iii) If $g=\mathrm{id}$, the statement follows from Proposition 4.1 as given the CY condition we have $\hat{c}=N-2\sum_{k}q_{k}=N-2$.
For $g\neq\mathrm{id}$, apply the same proposition again to estimate $\deg p$ in $\mathrm{Jac}(f^{g})$. Namely, it gives
@@
At the same time, we have $N_{g}+\mathrm{age}(g)=N-\mathrm{age}\bigl{(}g^{-1}\bigr{)}\leq N-1$ because $\mathrm{age}\bigl{(}g^{-1}\bigr{)}\in{\mathbb{N}}_{\geq 1}$. Combining this with the inequality above we get
$q_{l}(X)\leq N-2-\sum_{k\in I_{g}}q_{k}\leq N-2.$ (5.1)
@@
(iv) If $g=\mathrm{id}$, by Proposition 4.1, we see that $q_{l}([\mathrm{hess}(f)]\xi_{\mathrm{id}})=q_{r}([\mathrm{hess}(f)]\xi_{\mathrm{id}})=N-2$. If $g\neq\mathrm{id}$, $q_{l}(X)=q_{r}(X)=N-2$, then equation (5.1) implies that $\sum_{k\in I_{g}}q_{k}=0$ and, hence, $N_{g}=0$ and $\mathrm{age}(g)=\mathrm{age}\bigl{(}g^{-1}\bigr{)}=1$, which altogether implies
$N=\mathrm{age}(g)+\mathrm{age}\bigl{(}g^{-1}\bigr{)}+N_{g}=2.$
We now construct two symmetries of ${\mathcal{B}}_{\rm tot}$. The horizontal morphism $\Psi$ and the vertical morphism $\Phi$.
Consider the direct sum decomposition of ${\mathcal{B}}_{\rm tot}$ as in equation (4.1). We first extend the $\phi_{f,\beta}$ isomorphism of Proposition 4.2 to ${\mathcal{B}}_{\rm tot}$ in the following way. The hessian matrix of $f^{g}$ viewed coordinate free is a bilinear form on the tangent bundle of $\mathrm{Fix}(g)$. Therefore, its determinant $\mathrm{hess}(f^{g})$ is canonically an element of $\bigl{(}\Lambda^{N_{g}}\mathrm{Fix}(g)^{\vee}\bigr{)}^{\otimes 2}\otimes\mathbb{C}[\mathrm{Fix}(g)]$. Fix a generator
$\xi^{\vee}_{g}:=\iota_{\xi_{g}}{\rm d}x_{1}\wedge\cdots\wedge{\rm d}x_{N}\in\Lambda^{N_{g}}\mathrm{Fix}(g)^{\vee},$
where $\iota$ is the interior product operator and let the generator of $\bigl{(}\Lambda^{N_{g}}\mathrm{Fix}(g)^{\vee}\bigr{)}^{\otimes 2}$ to be $(\xi^{\vee}_{g})^{\otimes 2}$. This choice allows us to fix $\mathrm{hess}(f^{g})$ as a function on $\mathrm{Fix}(g)$ and, hence fix a pairing $\eta_{f^{g}}$ for the $g$-sector. As in Proposition 4.2, this in turn defines an isomorphism $\phi_{f^{g},\beta}$ on each sector. Now the vertical morphism $\Phi$ is the direct sum of these isomorphisms acting on each sector of ${\mathcal{B}}_{\rm tot}$
$\Phi:=\bigoplus_{g\in{\mathrm{GL}}_{f},\beta\in{\mathbb{Q}}}\phi_{f^{g},\beta}\colon\ {\mathcal{B}}_{\rm tot}\to{\mathcal{B}}_{\rm tot}^{\vee}.$
It is an isomorphism restricted to ${\mathcal{B}}_{{\rm tot},G}$ for any finite $G$ because each of $\phi_{f^{g},\beta}$ is an isomorphism.
Define the horizontal morphism $\Psi$ to act on the $g$-th sector by $\Psi([p]\xi_{g}):=[p]\xi_{g^{-1}}$. Extend it by linearity to all ${\mathcal{B}}_{\rm tot}$. This is an isomorphism because $f^{g}=f^{g^{-1}}$ and $\mathrm{Jac}(f^{g})=\mathrm{Jac}\bigl{(}f^{g^{-1}}\bigr{)}$.
Proposition 5.3.
  • $(1)$
    The maps $\Phi$ and $\Psi$ are well defined on ${\mathcal{B}}(f,G)$ for any finite $G\subseteq{\mathrm{SL}}_{f}$.
  • $(2)$
    For $f$ satisfying CY condition and a finite $G\subseteq{\mathrm{SL}}_{f}$, let $V^{a,b}$ stand for the bidegree $(a,b)$-subspace of ${\mathcal{B}}(f,G)$. Then the maps $\Psi$ and $\Phi$ induce the ${\mathbb{C}}$-vector spaces isomorphisms
[…]
@@
1) The map $\Psi$ commutes with the $G$-action since $\mathrm{Fix}(g)=\mathrm{Fix}\bigl{(}g^{-1}\bigr{)}$. Hence $\Psi$ preserves the invariants.
To see that $\Phi$ commute with the $G$-action, recall first that, $f\bigl{(}h^{-1}\cdot\mathbf{x}\bigr{)}=f(\mathbf{x})$ and, hence, $f^{g}\bigl{(}h^{-1}\cdot\mathbf{x}\bigr{)}=f^{hgh^{-1}}(\mathbf{x})$. Furthermore, since $\det(h)=1$ we have $\rho_{h,g}h^{*}\bigl{(}\xi_{hgh^{-1}}^{\vee}\bigr{)}=\xi^{\vee}_{g}$ and we can conclude that
$\rho_{h,g}^{2}\eta_{f^{hgh^{-1}}}\bigl{(}\bigl{[}p_{1}\bigl{(}h^{-1}\cdot\mathbf{x}\bigr{)}\bigr{]}\xi_{hgh^{-1}},\bigl{[}p_{2}\bigl{(}h^{-1}\cdot\mathbf{x}\bigr{)}\bigr{]}\xi_{hgh^{-1}}\bigr{)}=\eta_{f^{g}}([p_{1}(\mathbf{x})]\xi_{g},[p_{2}(\mathbf{x})]\xi_{g}).$
@@
To verify compatibility of $\Phi$ with the grading, note first, that $\widehat{c}(f^{g})=\sum_{k\in I_{g}}(1-2q_{k})$. Thus, by Proposition 4.2 the left charge of $[\phi_{f^{g},\deg p}(p)]\xi_{g}$ is given by
$q_{l}([\phi_{f^{g},\deg p}(p)]\xi_{g})$ $=\sum_{k\in I_{g}}(1-2q_{k})-\deg p-\sum_{k\in I_{g}^{c}}q_{k}+\mathrm{age}{(g)}$
$=N_{g}-2\sum_{k\in I_{g}}q_{k}-\deg p-\sum_{k\in I_{g}^{c}}q_{k}+\bigl{(}N-N_{g}-\mathrm{age}{\bigl{(}g^{-1}\bigr{)}}\bigr{)}$
$=-2+\sum_{k\in I^{c}_{g}}q_{k}-\deg p+N-\mathrm{age}{\bigl{(}g^{-1}\bigr{)}}=N-2-q_{r}([p]\xi_{g}).$
The computation for the right charge is identical. ∎
@@
Proposition 5.4.
For $f$ satisfying CY condition and a finite $G$, such that $J\subseteq G\subseteq{\mathrm{SL}}_{f}$, let $V^{a,b}$ stand for the bidegree $(a,b)$-subspace of ${\mathcal{B}}(f,G)$. Then
@@ 2. $(2)$
One notes immediately that $[1]\xi_{j_{f}}$ and $[1]\xi_{j_{f}^{-1}}$ are non-zero in ${\mathcal{B}}(f,G)$ and belong to $V^{0,N-2}$ and $V^{N-2,0}$ respectively. By Proposition 5.3 it is enough to show one of the statements.
Lemma 5.5.
$\dim V^{0,N-2}=\{g\in G\mid\mathrm{age}(g)=1$, $N_{g}=0\}$.
@@
Let $[p]\xi_{g}\in V^{0,N-2}$. It follows from equations (3.2) and (4.3) that $\mathrm{age}(g)=1-N_{g}/2$. The statement follows by Proposition 3.7. ∎
Under the CY condition for $g\in{\mathrm{GL}}_{f}\backslash\{\mathrm{id}\}$ of finite order and with integral $\mathrm{age}(g)$ we have by Proposition 3.14 that $\mathrm{age}(g)\geq 1$ with the equality being reached only for $g=j_{f}$. This completes the proof. ∎
This completes the proof of Theorem 1.1.
@@
$h^{a,b}:=\dim_{\mathbb{C}}\left\{X\in{\mathcal{B}}(f,G)\mid(q_{l}(X),q_{r}(X))=(a,b)\right\}$
@@
It follows from the propositions above that for $f$ satisfying CY condition and $G$ such that $J\subseteq G\subseteq{\mathrm{SL}}_{f}$, the numbers $h^{a,b}$ form a diamond,
[…]
Let us call the line $\bigl{\{}h^{a,b}\mid a+b=D\bigr{\}}$ the horizontal line and the line $\bigl{\{}h^{a,b}\mid a-b=0\bigr{\}}$ the vertical line. The Hodge diamond $\bigl{\{}h^{a,b}\bigr{\}}_{a,b=0}^{D}$ has the following special properties
@@
    The $g$-th sector of ${\mathcal{B}}(f,G)$ contributes as a line symmetric with respect to the horizontal line.
@@
    Every $g$-th sector of ${\mathcal{B}}(f,G)$ contributes together with a $g^{-1}$-th sector of ${\mathcal{B}}(f,G)$, located symmetrically with respect to the vertical line.
@@
$\xi_{j_{f}^{3}},\qquad\xi_{j_{f}^{5}},\qquad\xi_{j_{f}^{7}},\qquad\xi_{j_{f}^{9}},\qquad[x_{1}]\xi_{j_{f}^{4}},\qquad[x_{1}]\xi_{j_{f}^{8}},\qquad\bigl{[}x_{4}^{2}\bigr{]}\xi_{j_{f}^{6}},\qquad\bigl{[}x_{3}^{4}x_{4}^{4}\bigr{]}\xi_{\mathrm{id}},$
$\bigl{[}x_{1}x_{3}x_{4}^{4}\bigr{]}\xi_{\mathrm{id}},\qquad\bigl{[}x_{1}x_{3}^{3}x_{4}^{3}\bigr{]}\xi_{\mathrm{id}},\qquad\bigl{[}x_{2}x_{4}^{3}\bigr{]}\xi_{\mathrm{id}},\qquad\bigl{[}x_{1}^{2}x_{4}^{3}\bigr{]}\xi_{\mathrm{id}},\qquad\bigl{[}x_{1}x_{3}^{5}x_{4}^{2}\bigr{]}\xi_{\mathrm{id}},$
$\bigl{[}x_{2}x_{3}^{2}x_{4}^{2}\bigr{]}\xi_{\mathrm{id}},\qquad\bigl{[}x_{1}^{2}x_{3}^{2}x_{4}^{2}\bigr{]}\xi_{\mathrm{id}},\qquad\bigl{[}x_{2}x_{3}^{4}x_{4}\bigr{]}\xi_{\mathrm{id}},\qquad[x_{1}x_{2}x_{3}x_{4}]\xi_{\mathrm{id}},\qquad\bigl{[}x_{1}^{3}x_{3}x_{4}\bigr{]}\xi_{\mathrm{id}},$
$\bigl{[}x_{1}^{2}\bigr{]}\xi_{\mathrm{id}},\qquad\bigl{[}x_{2}^{2}\bigr{]}\xi_{\mathrm{id}}.$
@@
$\xi_{j_{f}},\qquad\xi_{j_{f}^{11}},\qquad[1]\xi_{\mathrm{id}},\qquad\bigl{[}x_{1}x_{2}^{2}x_{3}x_{4}^{4}\bigr{]}\xi_{\mathrm{id}},$
having the bigrading $(0,2)$, $(2,0)$, $(0,0)$ and $(2,2)$, respectively.
One gets the following diamond:
[…]
Example 5.7.
@@
$\xi_{\mathrm{id}}\qquad\text{and}\qquad\bigl{[}x_{1}^{3}x_{2}^{3}x_{3}^{3}x_{4}^{3}x_{5}^{3}\bigr{]}\xi_{\mathrm{id}}$
in bidegrees $(0,0)$ and $(3,3)$, respectively;
$[x_{1}x_{2}x_{3}x_{4}x_{5}]\xi_{\mathrm{id}},\qquad\bigl{[}x_{4}^{3}x_{5}^{2}\bigr{]}\xi_{\mathrm{id}},\qquad\bigl{[}x_{4}^{2}x_{5}^{3}\bigr{]}\xi_{\mathrm{id}},\qquad\bigl{[}x_{1}x_{2}x_{3}x_{4}^{2}\bigr{]}\xi_{\mathrm{id}},\qquad\bigl{[}x_{1}x_{2}x_{3}x_{5}^{2}\bigr{]}\xi_{\mathrm{id}},$
$\bigl{[}(x_{1}+x_{2}+x_{3})^{2}\bigr{]}(\xi_{(1,2,3)}+\xi_{(1,3,2)}),\qquad\bigl{[}x_{4}^{2}\bigr{]}(\xi_{(1,2,3)}+\xi_{(1,3,2)}),$
$\bigl{[}x_{5}^{2}\bigr{]}(\xi_{(1,2,3)}+\xi_{(1,3,2)}),\qquad[(x_{1}+x_{2}+x_{3})x_{4}](\xi_{(1,2,3)}+\xi_{(1,3,2)}),$
$[(x_{1}+x_{2}+x_{3})x_{5}](\xi_{(1,2,3)}+\xi_{(1,3,2)}),\qquad[x_{4}x_{5}](\xi_{(1,2,3)}+\xi_{(1,3,2)})$
@@
$\xi_{j_{f}}\qquad\text{and}\qquad\xi_{j_{f}^{4}}$
in bidegrees $(3,0)$ and $(0,3)$, respectively;
$\xi_{j_{f}^{2}},\qquad\xi_{(1,2,3)j_{f}}+\xi_{(1,3,2)j_{f}},\qquad\xi_{(1,2,3)j_{f}^{2}}+\xi_{(1,3,2)j_{f}^{2}}$